40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 23 21:30:29 2009, Jenne, Update, LSC, Measured MC length 
    Reply  Tue Nov 24 08:00:16 2009, rana, Update, LSC, Measured MC length 
       Reply  Tue Nov 24 10:36:21 2009, Koji, Update, LSC, Measured MC length 
          Reply  Wed Nov 25 03:05:15 2009, rob, Update, Locking, Measured MC length 
             Reply  Wed Nov 25 14:29:08 2009, rob, Update, Locking, Measured MC length--FSS trend FSStrendpowerjump.png
                Reply  Wed Nov 25 16:13:27 2009, rana, Update, PSL, Measured MC length--FSS trend 
                   Reply  Wed Nov 25 16:44:52 2009, Koji, Update, PSL, Measured MC length--FSS trend 091125_FSS.png
             Reply  Wed Nov 25 15:38:08 2009, rob, Update, Locking, Measured MC length 
Message ID: 2336     Entry time: Wed Nov 25 16:44:52 2009     In reply to: 2335
Author: Koji 
Type: Update 
Category: PSL 
Subject: Measured MC length--FSS trend 

I checked C1:PSL-FSS_VCODETPWR. The attached is the 4 months trend of the FSS RCTRANS / RFPDDC(=FSS REFL) / VCODETPWR / VCOMODLEVEL.

Although VCO modulation level setting was mostly constnt, VCODETPWR, which presumably represents the RF level, changes time by time.
It coincides with the recent reduction of the RCTRANS/RFPDDC. Actually, my touch restored the VCO to the previous more stable state.
One can see that this is not only a single occation, but it happened before too. (In the middle of Aug.)

This could be explained by the bad contact of some cable or connector.

Nevertheless we need more careful investigation:

1. Understand what VCODETPWR is exactly.
2. Investigate relationship between VCOMODLEVEL / VCODETPWR / AOM deflection efficiency / RCTRANSPD
3. Confirm the frequency matching between the VCO and AOM.

Quote:

but the increase in both the RCtrans and the RCrefl is consistent with my theory that the power going to the RC has increased ; its not just an increase in the visibility.

We should scan the AOM/VCO to make sure the frequency is matched to the resonance to within 0.5 dB.

 

Attachment 1: 091125_FSS.png  106 kB  | Hide | Hide all
091125_FSS.png
ELOG V3.1.3-